# Loss Landscape Degeneracy Drives Stagewise Development in Transformers

### v2

@@ F.2.1 Architecture
In the following $L$ refers to the number of layers (blocks) in the transformer, $H$ is the number of heads in each layer, $D$ is the dimension of inputs $x\in\mathbb{R}^{D}$ and $K$ is the number of $(x,y)$ pairs provided to the Transformer in-context.
The architecture is a pre-layer-norm decoder-only transformer modeled after NanoGPT Karpathy, 2022; see also Phuong and Hutter, 2022 with a learnable positional embedding. For the models discussed in the main body, we consider $L=2$, $H=4$ transformers (with $d=51,717$ parameters), i.e., two transformer blocks with four attention heads each.
@@
Token Embedding Weight $320$
Positional Embedding Weight $1,024$
Layer 1 Layer Norm Weight 2 $64$
Layer 1 Layer Norm Bias 1 $64$
Layer 1 Attention Weights $12,288$
Layer 1 Attention Output Weights $4,096$
Layer 1 Layer Norm Weight 1 $64$
Layer 1 Layer Norm Bias 2 $64$
Layer 1 Feed-Forward MLP Weight $4,096$
Layer 1 Feed-Forward MLP Bias $64$
Layer 1 Feed-Forward Output Weight $4,096$
Layer 1 Feed-Forward Output Bias $64$
Layer 2 Layer Norm Weight 1 $64$
Layer 2 Layer Norm Bias 1 $64$
Layer 2 Attention Weights $12,288$
Layer 2 Attention Output Weights $4,096$
Layer 2 Layer Norm Weight 2 $64$
Layer 2 Layer Norm Bias 2 $64$
Layer 2 Feed-Forward MLP Weight $4,096$
Layer 2 Feed-Forward MLP Bias $64$
Layer 2 Feed-Forward Output Weight $4,096$
Layer 2 Feed-Forward Output Bias $64$
Unembedding Layer Norm Weight 1 $64$
Unembedding Layer Norm Bias 1 $64$
Unembedding Weight 2 $320$
Unembedding Bias 2 $5$
Figure F.2: Transformer parameters in the in-context linear regression setting. The model has two transformer blocks for a total of $51,717$ trainable parameters.
@@ F.2.2 Tokenization
Through the main text, we write $f_{w}(S_{k})$ for $f_{w}(T_{k})$. Note that this tokenization includes the final $y_{k}$ token even though this receives no training signal. For this reason, we omit this token from the attention entropy and variability plots (Figures D.5 and D.6).
@@ F.2.3 Training
We train from a single seed for each choice of architecture and optimizer hyperparameters using minibatch stochastic gradient descent. We train without explicit regularization and use the Adam optimizer (Kingma and Ba, 2014). The training runs take 1 to 5 TPU-hours on TPUs provided by Google Research. Models are trained from the same initialization and on the data vectors within each batch (but for different sets of tasks and task orderings).

### v3

@@ F.2.1 Architecture
[…]
The architecture is a pre-layer-norm decoder-only transformer modeled after NanoGPT Karpathy, 2022; see also Phuong & Hutter, 2022 with a learnable positional embedding. For the models discussed in the main body, we consider $L=2$, $H=4$ transformers (with $d=51,717$ parameters), i.e., two transformer blocks with four attention heads each.
@@
Token Embedding Weight 320320320
Positional Embedding Weight 1,0241,0241 , 024
Layer 1 Layer Norm Weight 2 646464
Layer 1 Layer Norm Bias 1 646464
Layer 1 Attention Weights 12,28812,28812 , 288
Layer 1 Attention Output Weights 4,0964,0964 , 096
Layer 1 Layer Norm Weight 1 646464
Layer 1 Layer Norm Bias 2 646464
Layer 1 Feed-Forward MLP Weight 4,0964,0964 , 096
Layer 1 Feed-Forward MLP Bias 646464
Layer 1 Feed-Forward Output Weight 4,0964,0964 , 096
Layer 1 Feed-Forward Output Bias 646464
Layer 2 Layer Norm Weight 1 646464
Layer 2 Layer Norm Bias 1 646464
Layer 2 Attention Weights 12,28812,28812 , 288
Layer 2 Attention Output Weights 4,0964,0964 , 096
Layer 2 Layer Norm Weight 2 646464
Layer 2 Layer Norm Bias 2 646464
Layer 2 Feed-Forward MLP Weight 4,0964,0964 , 096
Layer 2 Feed-Forward MLP Bias 646464
Layer 2 Feed-Forward Output Weight 4,0964,0964 , 096
Layer 2 Feed-Forward Output Bias 646464
Unembedding Layer Norm Weight 1 646464
Unembedding Layer Norm Bias 1 646464
Unembedding Weight 2 320320320
Unembedding Bias 2 555
Figure F.2: Transformer parameters in the in-context linear regression setting. The model has two transformer blocks for a total of 51,71751,71751 , 717 trainable parameters.
@@ F.2.2 Tokenization
Through the main text, we write $f_{w}(S_{k})$ for $f_{w}(T_{k})$. Note that this tokenization includes the final $y_{k}$ token even though this receives no training signal. For this reason, we omit this token from the attention entropy and variability plots (Figures˜D.5 and D.6).
@@ F.2.3 Training
We train from a single seed for each choice of architecture and optimizer hyperparameters using minibatch stochastic gradient descent. We train without explicit regularization and use the Adam optimizer (Kingma & Ba, 2014). The training runs take 1 to 5 TPU-hours on TPUs provided by Google Research. Models are trained from the same initialization and on the data vectors within each batch (but for different sets of tasks and task orderings).